OpenStudy (anonymous):

Find the folowing limit: lim x->1 ((cos(x*pi/2))/(1-sqrt(x)))

hero (hero):

Hint: Multiply top and bottom by \(1 - \sqrt{x}\)

OpenStudy (anonymous):

you mean 1+sqrt(x)? im not in school, i have test today about this and this came up. so im interested how to solve it.

hero (hero):

Nope, I mean multiply top and bottom by \(1 - \sqrt{x}\). Multiplying an expression by 1 is legal. In this case, we just want to put it in a form where we can evaluate the limit right?

OpenStudy (anonymous):

i know that. but what to do when i get result of mutiplication?

hero (hero):

Because if you mulitplied \(\dfrac{\cos\left( \dfrac{\pi x}{2}\right)}{1 - \sqrt{x}} \dot\ \dfrac{1 - \sqrt{x}}{1 - \sqrt{x}}\), do you know the result of this multiplication?

OpenStudy (anonymous):

1-2sqrt(x)+x in denominator...

hero (hero):

Yes and from there, you should be able to evaluate the limit.

OpenStudy (anonymous):

no i am not :D thats why im asking ;)

hero (hero):

Hmmm. Interesting. It doesn't work that way either.

OpenStudy (anonymous):

i tried with 1+sqrt(x) and it fails to... result is pi, but im interested in how to get to that :D

hero (hero):

All you have to do is find a factor of one where upon multiplying it you get an expression in the denominator where x does not become zero when you evaluate the limit.

OpenStudy (anonymous):

yes, but its easy to say that :D

OpenStudy (anonymous):

Do you believe in magic? ^.^

OpenStudy (anonymous):

no :P

OpenStudy (anonymous):

Well you're about to :D Don't give up on multiplying by 1+sqrt(x)/1+sqrt(x) just yet. And you know... tell me what you get, because, well, I'm too lazy to LaTeX it out XD

OpenStudy (anonymous):

Oh wait... you haven't LaTeX'd anything at all... d'oy Fine, I'll do it.

OpenStudy (anonymous):

\[\Large \frac{(1+\sqrt x)\left[\cos\left(\frac{\pi}{2}x\right)\right]}{1-x}\]

OpenStudy (anonymous):

i can write in latex, but im too lazy to :) i got that...

OpenStudy (anonymous):

Okay. You can do it from here. Right? No? Haha... because you don't believe in magic >:)

OpenStudy (anonymous):

i cannot. :D

OpenStudy (anonymous):

if i decompose denominator im on start again :D

OpenStudy (anonymous):

And because I know you're smart, I'll let you work out the details of showing why \[\Large \cos\left[\frac\pi 2 - \frac\pi 2(1-x)\right]=\cos\left(\frac{\pi }{2}x\right)\] (Hint: It's just algebraic simplification)

OpenStudy (anonymous):

Now that THAT's out of the way, well, OBVIOUSLY \[\Large \cos\left[\frac\pi 2 - \frac\pi 2(1-x)\right]=\sin\left[\frac\pi2 (1-x)\right]\]

OpenStudy (anonymous):

ok i got it now

OpenStudy (anonymous):

So it all boils down to this: \[\Large \cos\left[\frac\pi2x\right]=\sin\left[\frac\pi2(1-x)\right]\] This matters? You bet it does ^.^

OpenStudy (anonymous):

i have sinx/x and thats trivial.... ty!!!

OpenStudy (anonymous):

Do you believe in magic? ^.^ Maybe not. But believe in me :D

OpenStudy (anonymous):

:D